Brief Exercise 15-8 Concord Inc. has outstanding 13,600 shares of $10 par value common stock. On July 1, 2017, Concord reacquired 105 shares at $87 per share. On September 1, Concord reissued 63 shares at $90 per share. On November 1, Concord reissued 42 shares at $85 per share. Prepare Concords journal entries to record these transactions using the cost method. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Debit Credit
Answer #1
7/1/17 Treasury stock 9135 =105*87
        Cash 9135
9/1/17 Cash 5670 =63*90
        Treasury Stock 5481 =63*87
        Paid-in Capital from Treasury Stock 189
11/1/17 Cash 3570 =42*85
Paid-in Capital from Treasury Stock 84
        Treasury Stock 3654 =42*87